I Need help on this one. Some corn costing 60 cents per kilogram is added to 50 kilograms of oats costing 90 cents per kilogram to make an animal feed costing 75 cents per kilogram. How many Kilograms of corn should be added.
So far i have .60(x+50)+.50(90)=.75+x

Answer by Paul(988)

60t+50(90-t)=90(75)
60t+4500-50t=6750 (subtract both sides by 3750 and collect like terms)
10t=2250
t=225
I am guessing 225kg of corn